OpenStudy (yacoub1993):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 7, b = 7, c = 5 A. A = 70°, B = 70°, C = 40° B. A = 69°, B = 69°, C = 42° C. A = 42°, B = 69°, C = 69° D. A = 69°, B = 42°, C = 69°

OpenStudy (yacoub1993):

I am sure it isnt D because i got wrong

OpenStudy (wolf1728):

The height or altitude of the triangle = hgt² = 7² - 2.5² hgt² = 49 -6.25 hgt² = 42.75 hgt = 6.538 angle A = arc tan (6.538 / 2.5) angle A = arc tan (2.6152) angle A = 69.0741 angle A = 69° angle B = angle A = 69° angle C = 180 -69 -69 angle C = 42° and there you have it !!!

OpenStudy (yacoub1993):

thnxs @wolf1728

OpenStudy (wolf1728):

This is the complete diagram of the triangle.

OpenStudy (wolf1728):

u r welcome yacoub

OpenStudy (yacoub1993):

i was confused with which law to solve for the first and second angles

OpenStudy (wolf1728):

I know that people here will mention about law of sines, law of cosines and so on. I think it is much easier to remember the trigonometric functions (sine = opp/hyp) and you can figure out everything else yourself.

OpenStudy (yacoub1993):

ohh ok i have another question which have no idea how to solve... Can u help me with it?

OpenStudy (wolf1728):

okay go ahead

OpenStudy (yacoub1993):

Find the quotient z1/z2 of the complex numbers. Leave answer in polar form. \[z1 = \frac{ 1 }{ 8 }(\cos \frac{ 2\pi }{ 3 }=i \sin \frac{ 2\pi }{ 3 })\] \[z2= \frac{ 1 }{ 3 }(\cos \frac{ \pi }{ 4 }+i \sin \frac{ \pi }{ 4 })\]

OpenStudy (yacoub1993):

@wolf1728 sory for late reply.....my internet conection went down